The perimeter of a rectangle is 100 cm and its diagonal is 32 cm find its area.
Answer by satyareddy22(84)

Let l be length,b be breadth
perimeter of a rectangle = 100 cm
l+b=50cm-----------(1)
diagonal = 32 cm
By Pythagorean theorem l^2+b^2=32^2=1024-----(2)
Squaring eq (1)
=>(l+b)^2=50^2
=>l^2+b^2 +2lb=2500
from eq (2)
1024+2lb=2500
2lb=2500-1024=1476
lb=1476/2=738
lb=738,which is area of rectangle(lb)